# LanPaint: Training-Free Diffusion Inpainting with Asymptotically Exact and Fast Conditional Sampling

### v2

@@ 1. The coefficients can take the following forms:
The hyperparameter $\alpha>0$ can be tuned based on the task. It represents the expected noise level of the sampling target, derived according to the forward diffusion process equation 2, under which a Gaussian random variable with standard deviation $\alpha$ follows the distribution $\mathcal{N}(0,1-\bar{\alpha}_{t}+\bar{\alpha}_{t}\alpha)$, whose score has linear term of the form $\frac{-1}{1-\bar{\alpha}_{t}+\bar{\alpha}_{t}\ \alpha}\mathbf{x}$. It is particularly useful when the assumption that $\hat{\mathbf{x}}_{0}(\mathbf{x})$ is constant does not hold. For instance, when sampling from a Gaussian distribution with standard deviation $\sigma$, $\hat{\mathbf{x}}_{0}(\mathbf{x})$ varies, and setting $\alpha=\sigma$ optimizes $A$. Thus, $\alpha$ can be interpreted as the "noise level" of the target distribution. For image generation tasks, set $\alpha=0$.
@@ 2. The coefficients can alternatively be expressed as:
This formulation incorporates the guidance scale $\lambda$ from equation 12, where $A$ scales proportionally with $\lambda$. The proportional relationship ensures stable solutions even at large $\lambda$ values. In practice, we adopt this set of parameter for the masked ($\mathbf{y}$, known) part of LanPaint.
Substituting these coefficients into the general solution (FLD Gaussian solution) yields exact analytical expressions for the FLD dynamics over an arbitrary time interval $[0,\tau]$. In practice, this can be adapted to a shifted interval $[\tau,\tau+\Delta\tau]$.
@@
Input: $\mathbf{x}_{0}$, $\mathbf{q}_{0}$, $\Delta\tau$, $\Gamma$, $A$, $D$, function $Ccoef$
@@
// Compute coefficients $\mathbf{C}$ via equation 82 or equation 83
@@
Input: $\mathbf{x}_{0}$, $\mathbf{q}_{0}$, $\Delta\tau$, $\Gamma$, $A$, $\mathbf{C}_{0}$, $D$, function $Ccoef$
@@
// Compute $\mathbf{C}$ via equation 82 or equation 83
@@
We employ the time step schedule inversely proportional to $A$
@@
The key insight comes from examining the exponential terms in equation 79. When $\tau=\Delta\tau$, most terms scale like $e^{-\Gamma\Delta\tau}$, where the product $\Gamma\Delta\tau$ directly determines the decay rate. By keeping $\Gamma\Delta\tau$ constant, we ensure a consistent "amount of change" per step—effectively balancing step size with the system’s intrinsic timescale. This approach automatically adjusts $\Delta\tau_{t}$ to be smaller when $\Gamma_{t}$ is large (fast dynamics) and larger when $\Gamma_{t}$ is small (slow dynamics), yielding stable and efficient numerical step across all diffusion time.
@@
$\begin{split}d\mathbf{x}&=\frac{1}{m}\mathbf{v}dt+\alpha\,\mathbf{s}(\mathbf{x})dt+\sqrt{2\alpha}dW_{\mathbf{x}}\\ d\mathbf{v}&=-\frac{\gamma}{m}\mathbf{v}dt+\,\mathbf{s}(\mathbf{x})dt+\sqrt{2\gamma}dW_{\mathbf{v}}\\ \end{split}$ (89)
@@
$\Psi=\alpha\gamma+1,\quad\mathbf{q}=\frac{\gamma}{\Psi}\frac{\mathbf{v}}{m},\quad\tau=\Psi\frac{t}{\gamma},\quad\Gamma=\frac{\gamma^{2}}{m\Psi},$
@@
$d\mathbf{x}=\mathbf{s}(\mathbf{x})\,d\tau+\sqrt{2}\,dW_{\tau},$
@@
$d\mathbf{x}=P\,\mathbf{s}(\mathbf{x})\,d\tau+\sqrt{2P}\,dW_{\tau},$
@@
[…]
@@
[…]
@@
[…]
@@
The FLD dynamics equation 65 can also be preconditioned by a positive definite symmetric matrix $P$, yielding:
@@
Appendix G Implementation
Mask Types
LanPaint
We implement LanPaint using the diffusers package, following Algorithm 4. Hyperparameters are configured as follows: $\gamma=15$, $\alpha=0.$, and $\lambda=8$ for image inpainting tasks on Celeb-A and ImageNet. The notation LanPaint-5 and LanPaint-10 denotes $N=5$ and $N=10$ sampling steps, respectively. The step size $\eta$ is set to $0.15$ for both Celeb-A and ImageNet. The impact of step size is ablated in Table.3, with the impact of other parameters discussed in Fig.7
Figure 7: Impact of expected noise $\alpha$, guidance scale $\lambda$, and friction $\gamma$ on LanPaint-10’s LPIPS for ImageNet box inpainting at stepsize 0.15, evaluated on a validation set of 100 images. Expected noise $\alpha$ most affects LPIPS, ideally 0 for image inpainting. Guidance scale $\lambda$ significantly improves performance from 0 (no guidance) to 4, with an optimal range of 6–10. Friction $\gamma$ has less significant effect, but small friction (too much momentum) reduces performance.
TFG
TFG (Cornwall et al. (2024)) corresponds to pure Langevin sampling. It is implemented via the Euler-Maruyama discretization, with step size schedule $\mathrm{d}\tau=\eta\sqrt{1-\bar{\alpha}_{t}}$ with $\eta=0.04$ as reported in their paper.
DDRM
Our implementation of DDRM follows Equations (7) and (8) in Kawar et al. Kawar et al. (2022), with $\sigma_{y}=0$, $\mathbf{V}=I$, and $s_{i}=1$ for observed regions $\mathbf{y}$ and $s_{i}=0$ for inpainted regions $\mathbf{x}$. The hyperparameters $\eta=0.7$ and $\eta_{b}=1$ are selected based on the optimal KID scores reported in Table 3 of the reference.
MCG
We implement MCG as described in Algorithm 1 of Chung et al. Chung et al. (2022b), using $\alpha=0.1/\|\mathbf{y}-P\hat{\mathbf{x}}_{0}\|$ as recommended in Appendix C. An alternative choice, $\alpha=0.1/\|\mathbf{y}-P\hat{\mathbf{x}}_{0}\|$, was evaluated but resulted in significantly poorer performance on ImageNet, as shown in Table 4. Consequently, we opted against using this alternative.
DPS
We adopt DPS Chung et al. (2022a) using $\alpha=1/\|\mathbf{y}-P\hat{\mathbf{x}}_{0}\|$ as recommended in Appendix D.
Alpha 0.1 Alpha 1.0
Method LPIPS FID LPIPS FID
CelebA-HQ-256
MCG 0.130 31.6 0.125 30.1
DPS 0.190 41.5 0.181 44.1
ImageNet
MCG 0.226 83.5 0.240 79.5
DPS 0.251 110.7 0.252 107.5
RePaint
We implement RePaint based on Algorithm 1 in Lugmayr et al. Lugmayr et al. (2022), with modifications to accommodate fast samplers. While the original method was designed for DDPM, we adapt it by replacing the backward sampling step (Line 7) with a Euler Discrete Sample step. Additionally, we set the jump step size to $1$ instead of the default $10$ backward steps recommended in Appendix B of the original work. This adjustment is necessary because fast samplers typically operate with only around $20$ backward sampling steps, making larger jump sizes impractical.
@@ 3. Fokker–Planck Equation and Stationary Distribution of Fast Langevin Dynamics
where $P$ is a symmetric positive semidefinite matrix. This system is the general form of underdamped Langevin dynamics, with preconditioning and the HFHR technique introduced in Appendix F. The Fokker–Planck equation for this system is:
@@
The FLD reparametrizes equation 110 through the transformations: $\mathbf{q}=\frac{\gamma}{m}\mathbf{v}$, $\tau=\frac{t}{\gamma}$, $\Gamma=\frac{\gamma^{2}}{m}$, $m=1$, $\alpha=0$, and $P=I$, resulting in the following system:
@@ Appendix I More Production-Level Model Evaluations Across Architectures
Figure 16: Model: hidream_i1_dev_fp8, Prompt: "a woman wearing a white lace camisole", Steps: 28, CFG Scale: 1.0, Sampler: Euler, Scheduler: Normal, LanPaint Iteration Steps: 5, Seed: 0, Batch Size: 4

### v3

@@ 1. The coefficients can take the following forms:
The hyperparameter $\alpha>0$ can be tuned based on the task. It represents the expected noise level of the sampling target, derived according to the forward diffusion process Eq.2, under which a Gaussian random variable with standard deviation $\alpha$ follows the distribution $\mathcal{N}(0,1-\bar{\alpha}_{t}+\bar{\alpha}_{t}\alpha)$, whose score has linear term of the form $\frac{-1}{1-\bar{\alpha}_{t}+\bar{\alpha}_{t}\ \alpha}\mathbf{x}$. It is particularly useful when the assumption that $\hat{\mathbf{x}}_{0}(\mathbf{x})$ is constant does not hold. For instance, when sampling from a Gaussian distribution with standard deviation $\sigma$, $\hat{\mathbf{x}}_{0}(\mathbf{x})$ varies, and setting $\alpha=\sigma$ optimizes $A$. Thus, $\alpha$ can be interpreted as the "noise level" of the target distribution. For image generation tasks, set $\alpha=0$.
@@ 2. The coefficients can alternatively be expressed as:
This formulation incorporates the guidance scale $\lambda$ from Eq.12, where $A$ scales proportionally with $\lambda$. The proportional relationship ensures stable solutions even at large $\lambda$ values. In practice, we adopt this set of parameter for the masked ($\mathbf{y}$, known) part of LanPaint.
Substituting these coefficients into the general solution (FLD Gaussian solution) yields exact analytical expressions for the FLD dynamics over an arbitrary time interval $0,\tau]$. In practice, this can be adapted to a shifted interval $\tau,\tau+\Delta\tau]$.
@@
Input: $\mathbf{x}_{0}$, $\mathbf{q}_{0}$, $\Delta\tau$, $\Gamma$, $A$, $D$, function CcoefCcoef
@@
// Compute coefficients $\mathbf{C}$ via Eq.121 or Eq.122
@@
Input: $\mathbf{x}_{0}$, $\mathbf{q}_{0}$, $\Delta\tau$, $\Gamma$, $A$, $\mathbf{C}_{0}$, $D$, function CcoefCcoef
@@
// Compute $\mathbf{C}$ via Eq.121 or Eq.122
@@
We employ the time step schedule inversely proportional to AA
@@
The key insight comes from examining the exponential terms in Eq.118. When $\tau=\Delta\tau$, most terms scale like $e^{-\Gamma\Delta\tau}$, where the product $\Gamma\Delta\tau$ directly determines the decay rate. By keeping $\Gamma\Delta\tau$ constant, we ensure a consistent "amount of change" per step—effectively balancing step size with the system’s intrinsic timescale. This approach automatically adjusts $\Delta\tau_{t}$ to be smaller when $\Gamma_{t}$ is large (fast dynamics) and larger when $\Gamma_{t}$ is small (slow dynamics), yielding stable and efficient numerical step across all diffusion time.
@@
[…]
@@
$\Psi=\alpha\gamma+1,\quad\mathbf{q}=\frac{\gamma}{\Psi}\frac{\mathbf{v}}{m},\quad\tau=\Psi\frac{t}{\gamma},\quad\Gamma=\frac{\gamma^{2}}{m\Psi},$ (129)
@@
$d\mathbf{x}=\mathbf{s}(\mathbf{x})\,d\tau+\sqrt{2}\,dW_{\tau},$ (132)
@@
$d\mathbf{x}=P\,\mathbf{s}(\mathbf{x})\,d\tau+\sqrt{2P}\,dW_{\tau},$ (133)
@@
$\pi^{\prime}(\mathbf{x}^{\prime})=\int\pi(\mathbf{x})\,\mathcal{N}(\mathbf{x}^{\prime};\mathbf{x}+P\mathbf{s}(\mathbf{x})\Delta\tau,2P\Delta\tau)\,d\mathbf{x},$ (134)
@@
$\pi^{\prime}(\mathbf{x}^{\prime})=\big(1-P\nabla\cdot\mathbf{s}(\mathbf{x}^{\prime})\Delta\tau\big)\int\pi(\mathbf{y}-P\mathbf{s}(\mathbf{x}^{\prime})\Delta\tau)\,\mathcal{N}(\mathbf{y};\mathbf{x}^{\prime},2P\Delta\tau)\,d\mathbf{y}.$ (138)
@@
$\int\pi(\mathbf{x}^{\prime}+\Delta\mathbf{x})\mathcal{N}\,d\mathbf{y}\approx\pi(\mathbf{x}^{\prime})-\Delta\tau\mathbf{s}(\mathbf{x}^{\prime})^{T}P\nabla\pi(\mathbf{x}^{\prime})+\Delta\tau P:\nabla\nabla\pi(\mathbf{x}^{\prime}).$ (141)
@@
The FLD dynamics Eq.101 can also be preconditioned by a positive definite symmetric matrix $P$, yielding:
@@ 3. Fokker–Planck Equation and Stationary Distribution of Fast Langevin Dynamics
where $P$ is a symmetric positive semidefinite matrix. This system is the general form of underdamped Langevin dynamics, with preconditioning and the HFHR technique introduced in Appendix G. The Fokker–Planck equation for this system is:
@@
The FLD reparametrizes Eq.152 through the transformations: $\mathbf{q}=\frac{\gamma}{m}\mathbf{v}$, $\tau=\frac{t}{\gamma}$, $\Gamma=\frac{\gamma^{2}}{m}$, $m=1$, $\alpha=0$, and $P=I$, resulting in the following system: